OpenStudy (anonymous):

What are two ways to simplify <the square root of 180.? You can show a square root by writing "sqrt()". For example, the square root of 2 can be written as "sqrt(2)

OpenStudy (anonymous):

mETAL & FAN

OpenStudy (anonymous):

Basic question is ;; what are two ways to simplify the square root of 180

OpenStudy (anonymous):

since \(180=36\times 5\) you have \[\sqrt{180}=\sqrt{36\times 5}=\sqrt{36}\times \sqrt5=6\sqrt5\]

OpenStudy (anonymous):

i don't know what "two ways" means

OpenStudy (imstuck):

maybe the positive and negative would be the 2?

OpenStudy (anonymous):

@kaylamarie081 Oh by the way it's 'Medal' not 'Metal'. Just thought you should know. Although some medals are made of metal, I think your going for awarding medals, not metals, to people who help you out. Although if it IS metals, and it's silver or gold, I'm on board.

OpenStudy (imstuck):

It's "you're", @Cosmichaotic, not "your".

OpenStudy (imstuck):

seeing as how we are pointing out spelling mistakes.

OpenStudy (anonymous):

Hahaha my bad, MEDAL not metal.

OpenStudy (anonymous):

Thank you @satellite73